A body weighing 64𝑙𝑏 slides down from rest at the top of a plane 18𝑓𝑑 long and inclined 30π‘œ above the horizontal. The coefficient of friction is 0.10. Find the velocity of the body as it reaches the bottom of the plane.


1
Expert's answer
2021-10-25T10:03:45-0400

The work-energy theorem gives:


"\frac{mv^2}{2}=W_1+W_2=-\mu mgl\cos\theta+mgl\sin\theta"

Hence, the speed of a body

"v=\sqrt{2gl(-\mu \cos\theta+\sin\theta)}\\
=\sqrt{2*32.2*18*(-0.10 \cos30^{\circ}+\sin30^{\circ})}\\
=22\:\rm ft/s"